a. An azimuth of 212 degrees is a bearing of what ? South _______A degrees West.

b. An oblique triangle has sides of 20 feet, 66 feet, 80 feet. What is the area of this triangle?

Compute to two decimal places
Answer #1

2 → 2190 KS 8 = 212-1860 320 The berring is 15 320W b) a= 25t, b=b.ft ca doft 5= a + b + c = 20 +6+80 = 83 ft Acea of terangl